find the equation of a circle with centre (-2 , 3) and point (1 , 2)

Answer by Boreal(15235)

The radius, r, is the distance between those two points. Distance formula is + sqrt (difference in x^2+difference in y^2)=sqrt(9+1)=sqrt(10)
for equation of circle, (x-h)^2+(y-k)^2=r^2, where (h, k) are coordinates of the center.
(x+2)^2+(y-3)^2=10